OpenStudy (anonymous):

-4y+8x=6 change this to slope intercept form(y=mx+b) identify m and b and then graph

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

y=2x-3/2

OpenStudy (anonymous):

4y=8x-6 y=2x-3/2

OpenStudy (anonymous):

adonis lopez...is that all right??

OpenStudy (anonymous):

for graphing u just find 2 points and plot. say x=0 , u get,y=-3/2 and y=1/2 u get x=1 so u hav points... (0,-3/2) and (1,1/2)